What is the empirical formula of a compound composed of 25.7 g potassium (K) and 5.26 g oxygen (O)? Ko Insert subscripts as needed.
Answer #1

Moles of K = 25.7/39 = 0.66 mole

Moles of O = 5.26/16 = 0.33 mole

Now, K:O = 0.66:0.33 = 2:1

So, empirical formula = K2O

If you have any query please comment

If you satisfied with the solution please rate it thanks...